OpenStudy (anonymous):

solve the system of equations by the substitution method y=4x+1 2y-4x=14

OpenStudy (anonymous):

2(4x+1) -4x = 14 8x - 4x + 2 = 14 4x = 12 x =3 y = 4*3 + 1 = 13 (3,13)

hero (hero):

y = 4x + 1 y = (14 + 4x)/2 2(4x+1) = 14 + 4x 8x + 2 = 14 + 4x 4x = 12 x = 3 y = 13